OpenStudy (anonymous):

a photograph is 6 in * 4 in. if you make a copy of the photograph with double the length and width, what is the area of the copy?

OpenStudy (anonymous):

length of new photo = 2*length of old photo = 2*6 = 12 width of new photo = 2*width of old photo = 2*4 = 8 area of new photo = l*w a = 12*8 a = 96in^2

OpenStudy (anonymous):

thank you